OpenStudy (anonymous):

Simply each expression :: -|-6-5|-13+(-5+15) is 12 :) Did i finally get it?

hartnn (hartnn):

whats |-6-5|= ?

OpenStudy (anonymous):

-1

hartnn (hartnn):

ummm...-6-5 = -11 |-11| = 11

OpenStudy (anonymous):

so its -6 + -5 not -6 - 5?

hartnn (hartnn):

-6 + -5 = -6 - 5 = -11

OpenStudy (anonymous):

so it is adding them? -2 -3 is -5 <<example

hartnn (hartnn):

yes. thats correct

hartnn (hartnn):

-2-3 = -(2+3) = -5

OpenStudy (anonymous):

so its -|-6-5|-13+(-5+15) -11 -13+10 -24 +10 -14

hartnn (hartnn):

that is so correct! good work :)

OpenStudy (anonymous):

ok thank you. let me see if i get the next question correct!

hartnn (hartnn):

welcome ^_^